OpenStudy (anonymous):

Calculus: how do you find Find dy/dx by implicit differentiation? : 3xy-x^2-4=0

OpenStudy (johnweldon1993):

Well you have a product there....so we can start by using the product rule on 3xy let f(x) = 3x and g(x) = y \[f'(x)g(x) + f(x)g'(x)\]

OpenStudy (johnweldon1993):

And remember that since you are differentiating with respect to x...you cannot differentiate 'y' ...it would simply become y'

OpenStudy (anonymous):

\[3xy-x ^{2}-4=0\] diff.w.r.t.x \[3\left( x \frac{ dy }{dx } +y*1\right)-2x=0,3x \frac{ dy }{dx }+3y=2x\] \[\frac{ dy }{ dx }=\frac{ 2x-3y }{ 3x }\]

OpenStudy (johnweldon1993):

so we have \[\frac{ d }{ dx } 3xy - x^2 - 4 = \frac{ d }{ dx }0\] via the product rule \[\frac{ d }{ dx } 3xy = 3\times y + 3x \times y'\] via the power rule \[\frac{ d }{ dx } -x^2 = -2x\] via the constant rule \[\frac{ d }{ dx } -4 = 0\] and again via the constant rule \[\frac{ d }{ dx } 0 = 0\] So altogether we have \[3y + 3xy' -2x = 0\] Add 2x and subtract 3y from both sides because we want to isolate the y' \[3xy' = -3y + 2x\] And finally to solve for y' we divide both sides by 3x \[y' = \frac{ -3y + 2x }{ 3x }\]

OpenStudy (johnweldon1993):

Good work @surjithayer ! medal for you!

OpenStudy (anonymous):

yw